OpenStudy (anonymous):

Divide: (x^6-7x^3+10)/(x^3-2)

jhonyy9 (jhonyy9):

x*6 -7x*3 +10 / x*3 -2 = x*3 - 5 x*6 -2x*3 --------- -- -5x*3 +10 -5x*3 +10 --------- -- -- - so this is : (x*3 -2)(x*3 -5)=x*6 -7x*3 +10

OpenStudy (anonymous):

is the stars powers

OpenStudy (anonymous):

http://goo.gl/FKDD2

OpenStudy (amistre64):

"shift+6" was just to far away :)